(Present value of an uneven stream of payments)

You are given three investment alternatives to analyze. The cash flows from these three investments are as follows:

a. What is the present value of investment A at an annual discount rate of 12  percent?

(Present value of an uneven stream of payments) You are given three investment alternatives to analyze The cash flows from these three investments are as follows Investment End of Year $ 1,000 2,000 3,000 (4,000) $2,000 2,000 2,000 2,000 4,000 $ 6,000 6,000 (6,000) (6,000) 4 4,000 16,000 hat is the present value of each of these three investments if the appropriate discount rate is 12 percent? a. What is the present value of investment A at an annual discount rate of 12 percent? S(Round to the nearest cent.)
Answer #1

Present value of investment A=1000/(1+12%)^1+2000/(1+12%)^2+3000/(1+12%)^3-4000/(1+12%)^4+4000/(1+12%)^5

=4350.22

Present value of investment B=2000/(1+12%)^1+2000/(1+12%)^2+2000/(1+12%)^3+2000/(1+12%)^4+4000/(1+12%)^5

=8344.41

Present value of investment C=6000/(1+12%)^1+6000/(1+12%)^2-6000/(1+12%)^3-6000/(1+12%)^4+16000/(1+12%)^5

=11135.35


answered by: TTT